• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Jon Miller

    Log entry time 08:01:35 on March 6,2007

    Entry number 194925

    keyword=Shift Summary

    Jon Miller
    Byung Lee

    00:25 MCC informed us that the beam would trip at 50 uA. So running ~35 uA instead.
    00:20 Moved target to Carbon
    00:37 MCC informed us that they could move the current up to 50 uA.
    01:43 Spectrometer Move completed
    01:52 Raster is ON
    02:46 MCC called and said that they were going to tune the beam to try to find the problems that was causing the recent beam trips.
    03:07 Raster is turned Off, beam is 50 uA.
    04:11 Asked for mask for target change to Pb.

    1194 00:23 Carbon Kin01
    1195 01:51 Carbon Kin02 Raster ON
    1196 03:11 Carbon Kin02 Raster Off
    1197 04:25 Lead Kin02 Raster ON
    1198 05;24 Lead Kin02
    1199 06:22 Lead Kin02
    1200 07:18 Lead Kin02